# Parameter-Efficient Fine-Tuning with Attributed Patch Semantic Graph for Automated Patch Correctness Assessment

### v1

@@ 6 Experimental Result
Our 77.2% 78.8% 79.3% 79.0%
TABLE XII: Effectiveness of Graph-LoRA in a cross-project setting on the Lin dataset.
@@
Our 87.9% 88.2% 88.0% 88.1%
@@
Through the above experiments, we have demonstrated that the performance of our method for the APCA task is optimal in a cross-validation setting. However, in practical applications, model needs to assess patches from unseen projects. To further explore the performance of our method, we design a cross-project verification using the Wang, Merge, Balance, and Lin datasets. For example, we use patches from projects other than Chart to train the model and use patches from Chart to evaluate the model. As the Merge dataset and Balance dataset contain patch data from projects other than the five listed projects Chart, Closure, Lang, Math, and Time, we regard these patch data from other projects as training data. In the experiments, we use three state-of-the-art APCA methods as baseline models, including CACHE, APPT, and LLM4PatchCorrect. To effectively evaluate the performance upper limit of the APCA models, our model and the LM4PatchCorrect model are both based on the LLM Llama3.
@@
Table IX, Table X, Table XI, and Table XII show the results of cross-project prediction on the Wang, Merge, Balance, and Lin datasets respectively. From the tables, we can see that in the cross-project prediction scenario, the accuracy, precision, recall, and F1 score of our model are 91.7%, 92.3%, 92.5%, 92.4% on the Wang dataset, 92.8%, 93.4%, 92.7%, 93.0% on the Merge dataset, 77.2%, 78.8%, 79.3%, 79.0% on the Balance dataset, and 87.9%, 88.2%, 88.0% and 88.1% on the Lin dataset. From the results, we can see that the performance of our model has declined in processing unseen patches. However, note that our model still outperforms all APCA baseline models. Compared with the best model APPT (based on traditional machine learning), our model has improved accuracy, precision, recall, and F1 score by 7.6% to 15.4%, 8.1% to 13.1%, 7.0% to 10.4%, and 7.5% to 14.0% respectively. Compared with the best model LLM4PatchCorrect (based on LLM), our model has improved accuracy, precision, recall, and F1 score by 1.6% to 3.6%, 1.8% to 3.7%, 0.9% to 3.6%, and 1.4% to 3.6% respectively.
Answer to RQ3: The performance under a cross-project scenario demonstrates that: (1) Compared with the cross-validation setting, the performance of APCA models in the cross-project scenario generally deteriorates; (2) In the cross-project scenario, our model still achieves the state-of-the-art performance using all metrics and datasets.
@@ 7.2 Threats to Validity
Threats to external validity. A threat to external validity is related with whether our results can be generalized. To minimize this threat, 1) we conduct experiments on four widely used APCA datasets, ranging from big to medium to small in terms of size and ranging from balanced to imbalanced in terms of the ratio between the correct and overfitting patches; 2) we also consider three different representative LLMs when LLM is involved with in this study; 3) with regard to the baselines, we always select the most representative and state-of-the-art techniques in each category of existing works on APCA task. Another threat to external validity is related with the implementation. Our implementation currently supports Java language only, and further efforts are needed to apply our approach to other programming languages. We consider addressing this limitation as an important direction for future work.
Threats to internal validity. One threat to internal validity is that we can possibly introduce errors during the experimental process. To reduce this threat as much as possible, several authors have carefully and independently examined the artifacts. Besides, to facilitate the replication and verification of our work, we have made the relevant materials (including code, datasets, models, etc.) publicly available for the community to review. Another threat to internal validity concerns the use of LLM, and the issue is that the LLM during the pre-training process may possibly have encountered the content of the used datasets. However, this is a common potential issue faced by most studies that use LLMs for code related tasks. In particular, note that this potential issue is also faced by the baseline method LLM4PatchCorrect in our experiment. Using the same LLM, our method consistently demonstrates clear advantages over LLM4PatchCorrect. This suggests that our method itself offers new insights for statically predicting patch correctness.
@@ 8 Conclusion
Patch overfitting is a serious issue which overshadows the automated program repair area, and many research efforts have been devoted for automated patch correctness assessment (APCA). With the emergence of large language model (LLM) technology, researchers have employed LLM to assess the patch correctness. The literature on APCA has highlighted the importance of capturing patch semantic and explicit code attributes in predicting patch correctness. However, existing LLM-based methods 1) typically treat code as token sequences and ignore the inherent formal structure for code, and 2) do not explicitly account for enough code attributes. To overcome these drawbacks, we in this paper design a novel patch graph representation named attributed patch semantic graph (APSG), which adequately captures the patch semantic and explicitly reflects important patch attributes. To effectively use graph information in APSG, we accordingly propose a new parameter-efficient fine-tuning (PEFT) method of LLMs named Graph-LoRA. The results of extensive evaluations show that compared to the state-of-the-art methods, our method improves accuracy and F1 score by 2.3% to 6.6% and 1.8% to 6.1% respectively. For future work, we will apply our method to more LLMs and demonstrate the effectiveness of our method in more code related tasks.

### v2

@@ 6 Experimental Result
Our 79.6% 81.6% 82.3% 81.9%
TABLE XVI: Effectiveness of our method in a cross-project setting on the Lin dataset.
@@
Our 88.5% 91.6% 90.3% 90.9%
TABLE XVII: Effectiveness of our method in a cross-project setting on the Multi-Benchmarks dataset.
Project Approach Accuracy Precision Recall F1
Chart APPT 77.5% 75.0% 78.4% 76.7%
LLM4PatchCorrect 89.8% 90.5% 90.7% 90.6%
Our 92.3% 92.4% 92.8% 92.6%
Closure APPT 70.9% 73.3% 75.8% 76.9%
LLM4PatchCorrect 82.4% 85.6% 86.2% 85.9%
Our 86.9% 89.1% 89.8% 89.4%
Lang APPT 77.3% 77.6% 75.0% 76.3%
LLM4PatchCorrect 85.8% 85.2% 87.3% 86.2%
Our 90.6% 90.2% 91.1% 90.6%
Math APPT 73.1% 74.5% 78.3% 76.4%
LLM4PatchCorrect 87.2% 88.9% 89.8% 89.3%
Our 91.1% 91.3% 91.8% 91.5%
Time APPT 84.0% 75.4% 70.7% 73.0%
LLM4PatchCorrect 90.5% 91.6% 91.3% 91.4%
Our 92.3% 93.1% 93.4% 93.2%
Average APPT 76.6% 75.2% 75.6% 75.9%
LLM4PatchCorrect 87.1% 88.4% 89.1% 88.7%
Our 90.6% 91.2% 91.8% 91.5%
@@
Through the above experiments, we have demonstrated that the performance of our method for the APCA task is optimal in a cross-validation setting. However, in practical applications, the model needs to assess patches from unseen projects. To further explore the performance of our method, we conduct the cross-project verification using the Wang, Merge, Balance, Lin, and Multi-Benchmarks datasets. For example, with regard to a certain dataset, we use patches from projects other than Chart to train the model and use patches from Chart to evaluate the model. As the three datasets Merge, Balance, and Multi-Benchmarks contain patch data from projects other than the five listed projects Chart, Closure, Lang, Math, and Time, we regard these patch data from other projects as training data when conducting cross-project verification on these three datasets. In the experiments, we use two state-of-the-art APCA methods as baseline methods, including APPT and LLM4PatchCorrect. To effectively evaluate the performance upper limit of the APCA methods, our method and the LM4PatchCorrect method are both based on the LLM Llama3.
Table XIII, Table XIV, Table XV, Table XVI, and Table XVII show the results of cross-project prediction on the Wang, Merge, Balance, Lin, and Multi-Benchmarks datasets respectively. From the tables, we can see that in the cross-project prediction scenario, the accuracy, precision, recall, and F1 score of our method are 92.4%, 93.1%, 93.3%, 93.2% respectively on the Wang dataset, 93.4%, 94.0%, 93.4%, 93.7% respectively on the Merge dataset, 79.6%, 81.6%, 82.3%, 81.9% respectively on the Balance dataset, 88.5% 91.6% 90.3% 90.9% respectively on the Lin dataset, and 90.6%, 91.2%, 91.8%, 91.5% respectively on the Multi-Benchmarks dataset. From the results, we can see that the performance of our method has declined in processing unseen patches. However, note that our method still outperforms all APCA baseline models. Compared with the best APPT method (the state-of-the-art ML-based method), our method has improved accuracy, precision, recall, and F1 score by 8.2% to 16.0%, 8.7% to 20.8%, 7.7% to 16.2%, and 8.2% to 16.8% respectively. Compared with the LLM4PatchCorrect methdo (the most advanced LLM-based method), our method has improved accuracy, precision, recall, and F1 score by 2.2% to 6.0%, 2.8% to 6.5%, 2.1% to 6.6%, and 2.3% to 6.5% respectively.
Answer to RQ3: The performance under a cross-project scenario demonstrates that: (1) Compared with the cross-validation setting, the performance of APCA methods in the cross-project scenario generally deteriorates; (2) In the cross-project scenario, our method still achieves the state-of-the-art performance using all metrics and datasets.
@@ 7.1 Case Study
7.2 Effect of the Ground-Truth Patches
In this section, we aim to explore the effect of ground-truth patches on the model performance. In the APCA task, ground-truth patches are typically used for APR experimentation [yang2023large, overfittingsurvey], where ground-truth patches may be accessible and plausible patches that behave differently from the ground-truth patches are deemed as overfitting. However, note that the ground-truth patch information is unavailable for real-world bug fixing, and our method focuses on this scenario and does not rely on ground-truth data. In fact, there are already some APCA works that use ground-truth patches to assess patch correctness. For example, Xin et al. [xin2017identifying] propose DiffTGen, which generates new tests based on the execution process of the ground-truth patches to improve the robustness of the method. Ye et al. [ye2021scale] compare the differences in runtime information between ground-truth patches and APR-generated patches to assess patch correctness.
To investigate the effect of ground-truth patches on the model performance, we use them as prompt inputs to the model and conduct experiments on the Lin dataset and the Balance dataset. More specifically, we first collect Defects4J human-written patches as ground-truth patches. Then, for each patch to be assessed, we construct a patch pair consisting of the patch itself and the ground-truth patch for the corresponding bug. For example, in the Lin dataset, there are 17 patches to be assessed for Chart-1. We thus construct 17 patch pairs using each patch to be assessed and the ground-truth patch for the corresponding bug. Next, during data preprocessing (i.e., the lower half of step (a) in Fig. 2), we place the ground-truth patch in the patch pair (preceding the patch to be assessed) as a prompt to feed into the model. The input with prompt of ground-truth patch is shown in Fig. 5. Finally, we train and test the model with ground-truth prompt in the same experimental settings as RQ1. Note that when evaluating the performance of the model with prompt of ground-truth patch, we remove the human-written patches from the evaluation data to prevent data leakage. The results are given in Table XVIII.
Figure 5: The input with prompt of ground-truth patch.
From the table, we can see that in the Balance dataset, the performance of the model with prompt of ground-truth patch improves by 1.5%, 1.3%, 1.1%, and 1.2% in terms of the accuracy, precision, recall, and F1 score respectively. In the Lin dataset, the performance of the model with prompt of ground-truth patch improves by 0.7%, 0.8%, 0.6%, and 0.7% in terms of the accuracy, precision, recall, and F1 score respectively. The results suggest that after adding prompt of ground-truth patch, the performance of the model improves by varying degrees on the two datasets. The improvement is more remarkable on the Balance dataset, which contains fewer patch samples. Overall, we believe that ground-truth patches can help the model assess patch correctness, especially when the data is scarce.
TABLE XVIII: The result of adding the prompt of ground-truth patch to the model.
Dataset Model Accuracy Precision Recall F1
Balance dataset Graph-LoRA-Llama3 86.7% 87.8% 87.2% 87.5%
Graph-LoRA-Llama3
+Ground-Truth
88.2% 89.1% 88.3% 88.7%
Lin dataset Graph-LoRA-Llama3 91.8% 92.5% 92.2% 92.3%
Graph-LoRA-Llama3
+Ground-Truth
92.5% 93.3% 92.8% 93.0%
@@ 7.3 Threats to Validity
Threats to external validity. A threat to external validity is related with whether our results can be generalized. To minimize this threat, 1) we conduct experiments on five APCA datasets, ranging in size from large to small, vary from one to multiple in terms of the number of bug benchmarks used for constructing the datasets, and vary from balanced to imbalanced in terms of the ratio between the number of correct patches and the number of overfitting patches; 2) we also consider three different representative LLMs when LLM is involved with in this study; 3) with regard to the baselines, we always select the most representative and state-of-the-art techniques in each category of existing works on APCA task. Another threat to external validity is related with the implementation. Our implementation currently supports Java language only, and further efforts are needed to apply our approach to other programming languages. We consider addressing this limitation as an important direction for future work.
Threats to internal validity. One threat to internal validity is that we can possibly introduce errors during the experimental process. To reduce this threat as much as possible, several authors have carefully and independently examined the artifacts. Besides, to facilitate the replication and verification of our work, we have made the relevant materials (including code, datasets, models, etc.) publicly available for the community to review. Another threat to internal validity concerns the use of LLM, and the issue is that the LLM during the pre-training process may possibly have encountered the content of the used datasets. However, this is a common potential issue faced by most studies that use LLMs for code related tasks. In particular, note that this potential issue is also faced by the baseline method LLM4PatchCorrect in our experiment. Using the same LLM, our method consistently demonstrates clear advantages over LLM4PatchCorrect. This suggests that our method itself offers new insights for statically predicting patch correctness. Meanwhile, the results of the ablation study (specifically the part of discarding training and only giving simple prompt) in Section 6.2 also suggest that the three LLMs we used likely have no data leakage issue on the data datasets we used.
@@ 8 Conclusion
Patch overfitting is a serious issue which overshadows the automated program repair area, and many research efforts have been devoted for automated patch correctness assessment (APCA). With the emergence of large language model (LLM) technology, researchers have employed LLM to assess the patch correctness. The literature on APCA has highlighted the importance of capturing patch semantic and explicit code attributes in predicting patch correctness. However, existing LLM-based methods 1) typically treat code as token sequences and ignore the inherent formal structure for code, and 2) do not explicitly account for enough code attributes. To overcome these drawbacks, we in this paper design a novel patch graph representation named attributed patch semantic graph (APSG), which adequately captures the patch semantic and explicitly reflects important patch attributes. To effectively use graph information in APSG, we accordingly propose a new parameter-efficient fine-tuning (PEFT) method of LLMs named Graph-LoRA. The results of extensive evaluations show that compared to the state-of-the-art methods, our method improves the accuracy and F1 score by 3.1% to 7.5% and 3.0% to 7.1% respectively. For future work, we will apply our method to more LLMs and demonstrate the effectiveness of our method in more code related tasks.
Acknowledgments
We deeply appreciate the anonymous reviewers for their insightful comments. This work was supported by National Natural Science Foundation of China (Grant No. 62102233), Shandong Province Overseas Outstanding Youth Fund (Grant No. 2022HWYQ-043), Joint Key Funds of National Natural Science Foundation of China (Grant No. U24A20244), and Qilu Young Scholar Program of Shandong University.